OpenStudy (anonymous):

x,y€ R if xy^2+3yx^2=5, then the value maximum of x+y is...?

OpenStudy (anonymous):

\[xy^2+3x^2y=5\] \[y^2+3xy=5/x\] \[y+3x=5/xy\] \[y+x = 5/xy - 2x\] lemme see if i can get an interval

OpenStudy (anonymous):

do you know the basic conditions for maxima and minima?

OpenStudy (anonymous):

help.. how to solve this problem, using differential ta??